# SoK: What don’t we know? Understanding Security Vulnerabilities in SNARKs

### v2

@@ 1 Introduction
On the forefront of the practical application of general-purpose ZKPs are Succinct Non-interactive Argument of Knowledge (SNARKs) [47, 82, 52, 43, 25]. SNARKs are non-interactive protocols that allow the prover to generate a succinct proof. The proof is efficiently checked by the verifier, while maintaining three crucial properties: completeness, soundness, and zero-knowledge. What makes SNARKs particularly appealing is their general-purpose nature, allowing any computational statement represented as a circuit to be proven and efficiently verified. Typically, SNARKs are used to prove that for a given function $f$ and a public input $x$, the prover knows a (private) witness $w$, such as $f(x,w)=y$. This capability allows SNARKs to be used in various applications, including ensuring data storage integrity [89], enhancing privacy in digital asset transfers [69, 93] and program execution [16, 14], as well as scaling blockchain infrastructure [86, 85, 62, 96]. Their versatility also extends to non-blockchain uses, such as in secure communication protocols [92, 107, 64] and in efforts to combat disinformation [57, 31, 59]. Unfortunately, developing and deploying systems that use SNARKs safely is a challenging task.
@@ 2 Background on SNARKs
[…]
@@ 3.1 System Model
Circuit Embedded Domain Specific Languages. An Embedded Domain Specific Language (eDSL) is a type of domain-specific language that is embedded within a host general-purpose programming language. For developing SNARKs, several eDSLs have emerged in recent years, embedded as libraries in Golang [15], Rust [106, 36, 3, 88, 109], TypeScript [73], C++ [94] and Java [60]. eDSLs can seamlessly interact with other code written in the host language, allowing easy integration with existing libraries. At the same time, they require developers to actively distinguish between in-circuit and out-of-circuit operations, which requires domain expertise to ensure correct implementation.
@@
Circuit Compiler. The circuit compiler in a SNARK implementation compiles the circuit specified in a high-level DSL or eDSL to the respective arithmetization. Note, that not all compilers target the compilation of high-level languages to specific arithmetizations. For example, high-level programs can first be compiled to VampIR [2] and ACIR [72], which are intermediate representations that allow compilation to, e.g., R1CS or some form of Plonkish arithmetization and aim to ease the support of multiple backends. Similarly, zkLLVM provides a compiler from the LLVM IR to different arithmetizations, allowing users to prove the execution of, e.g., native Rust, Golang, or C++ [40].
@@
Setup. In pre-processing SNARKs, the setup algorithm generates the prover and verifier keys ($pk$ and $vk$) by encoding the circuit relation, with the primary goal of succinct verification and optimized proving. However, different SNARKs have vastly differing properties with regard to trust assumptions, which in turn also impact their performance. Some SNARKs require a trusted setup, i.e., a randomized setup phase that introduces a trapdoor either per circuit (i.e., non-universal) [52] or in a universal setup phase [43]. The randomized part of the setup involves public parameters and random input to generate a common reference string (CRS) [71], which can be used to derive $pk$ and $vk$ in the deterministic part of the setup for a concrete circuit instantiation. The trapdoor needs to be discarded after the setup proccess is done, as anyone who has knowledge of the trapdoor can forge proofs. SNARKs that do not employ a trapdoor in the setup phase are commonly coined transparent [106, 12]. Note, that this introduces a trade-off in verifier efficiency — transparent SNARKs only have poly-logarithmic-time verification as compared to constant-time for SNARKs with trusted setup [99].
@@
Prover & Verifier. The prover consumes the values generated by the frontend, where the circuit is represented in a specific arithmetization. At a high level, the prover of a pre-processing SNARK (aside from Groth16 [52]) utilizes a polynomial commitment to commit to the satisfying witness, and successively engages with the verifier in an interactive protocol to evaluate the correctness of assignments with regard to the circuit relation. This “recipe” leads to a SNARK if the verifier is public-coin [83], and the prover can hash the transcript of the interactive proof to render the protocol non-interactive (a process commonly denoted as applying the “Fiat-Shamir heuristic”). Different proof systems choose different combinations of polynomial commitments and interactive proofs, resulting in differing properties for the respective target applications. The most popular approaches include Plonk [43], which employs a constant-round polynomial interactive oracle proof (IOP) with the KZG commitment [58]. Halo2 [106] uses the Plonk IOP with the Bulletproofs polynomial commitment[18]. Another class of protocol combines a polynomial commitment base on Merkle hashes with a low-degree test, resulting in a protocol that does not demand for a trusted setup or operations in elliptic curve groups [11], at the cost of non-constant verification and a larger proof size [34]. Plonky2 [88] uses the FRI-based polynomial commitment scheme and combines it with the Plonk IOP. Groth16 [52], the SNARK with the shortest proof size, is based on probabilistically checkable proofs (PCP) [47] and pairings over elliptic curves for proof verification. For an in-depth taxonomy, we refer readers to [99]. Some applications leverage one of the above SNARKs without the zero-knowledge property, solely relying on non-interactivity and succinctness (e.g zk-EVMs [96, 86]). In most SNARKs, zero-knowledge can be obtained cheaply with minor modifications.
@@ 4 Methodology
Analyzed SNARK Implementations. Our examination focuses on widely deployed SNARK systems, including ZK-rollups for blockchain scalability (e.g., zkSync Era [62], Polygon ZK-EVM [86] Scroll [96]), [^1] Notably, ZK-rollups have more than 1B USD in accumulated value locked in them according to L2BEAT https://l2beat.com/scaling/summary privacy-centric blockchains (e.g., Zcash [93], Aztec [6], Aleo [39]), and decentralized privacy applications (e.g., Tornado Cash [84]). We also assess key circuit libraries and SNARK frameworks (e.g., Circom and halo2) [34]. Our analysis further extends to inspecting the underlying proof systems for known bugs. This comprehensive approach ensures that we cover a broad spectrum of real-world SNARK applications and the various layers of technology they rely on.
Data Sources. In our comprehensive study of security vulnerabilities related to SNARK systems, we have utilized an extensive range of data sources to ensure a thorough analysis. Our primary source includes critical and high-impact vulnerabilities identified in $107$ security audit reports of systems employing SNARKs. Among these, $75$ reports specifically highlighted vulnerabilities directly related to SNARKs, while the remaining reports detailed other types of vulnerabilities, such as common smart contract deficits. Additionally, we have incorporated vulnerabilities from public bug bounty programs and vulnerability disclosures. These sources provide real-world examples of how vulnerabilities have impacted operational systems as well as the measures taken to address them. Furthermore, we delved into bug trackers of popular GitHub projects related to SNARKs. Our focus here has been specifically on security-related bugs, filtering out non-security-related issues. Table 2 in our paper depicts these varied data sources, collectively forming the foundation of our study of SNARK security.
@@ 8 Defenses For SNARKs
Despite these initial strides in tool development for ZK circuits, significant limitations remain. Tools relying on SMT solvers, like Picus and Korrect, face challenges due to limited support and efficiency in handling finite field arithmetic, leading to performance bottlenecks [54, 78]. Static analysis tools are often restricted in the range of vulnerabilities they can detect and tend to be specific to certain languages. CirC [77] compiler represents a step towards language-agnostic tooling, allowing various DSLs to compile into an Intermediate Representation, which can then be analyzed for potential vulnerabilities. Another promising direction is the creation of more secure DSLs, like Aleo [27], noname [109], and Noir [72], or eDSLs such as o1js [73] and arkworks R1CS [3], which incorporate strong typing and improved abstractions to prevent common issues.
@@
Multi-Provers as a Defense Mechanism. SNARK proofs are often used in safety-critical contexts, yet, with the current state of technology, guaranteeing that a complex SNARK system is bug-free and safe remains an elusive goal. The multi-prover design enhances safety by introducing redundancy at the proof system level. A multi-prover system utilizes multiple proof systems besides the primary SNARK, such as alternative SNARK proofs or TEE-based proofs, and primary proofs are accepted only if the secondary ones agree with it. This design trades off liveness with more safety, as a dispute will potentially bring the system to a halt. Such systems require a dispute resolution mechanism to handle potential disagreements. In practice, some ZK-Rollup projects [96], arguably one of the most complex categories of SNARK systems today, are already experimenting with multi-provers. [^5] https://scroll.io/blog/scaling-security
@@ 10 Discussion
2) Insight – Soundness bugs affecting SNARK verifiers lead on average to high severity bugs. Particularly when soundness bugs are exploitable, typically due to under-constrained circuits, they can lead to significant security breaches. Their potential to compromise system security has led to suboptimal strategies, such as employing multi-provers and permissioned provers, alongside significant bug bounties, sometimes reaching up to \$500k for a single vulnerability. These measures underscore the critical nature of soundness in maintaining the trustworthiness of SNARK-based systems.
3) Insight – Low-level circuit DSLs are easy-to-misuse leading to many vulnerabilities. Crafting efficient ZK circuits often necessitates using low-level ZK DSLs such as Circom and Gnark, reminiscent of early assembly and C programming, where common high-level programming features such as abstractions and basic types are absent. This complexity not only steepens the learning curve for developers but also increases the likelihood of introducing vulnerabilities into the circuits. The prevalent use of tools such tools highlights a regression to an era with bugs due to the absence of modern programming safeguards, underscoring an urgent need for more user-friendly DSLs that offer better abstractions and safety features, thereby mitigating the main shortcomings and reducing the vulnerability surface of SNARK applications.
@@ Acknowledgments
We thank Shankara Pailoor, Alex Ozdemir, Jean-Philippe Aumasson, and Justin Traglia for their insightful feedback. We thank the Ethereum Foundation and 0xPARC for supporting this work.
@@ A.2 Frozen Heart – Insecure Fiat Shamir Transformation
Note that if at any point a malicious prover does not like the challenge produced by Fiat-Shamir, his only option is to change his previous messages, but given that the hash function is cryptographically secure, he is unable to efficiently find a suitable change. However, invoking Fiat-Shamir with only a partial transcript of the protocol can create serious soundness issues, referred to as Frozen Heart attacks [75]. Intuitively, this enables a malicious prover to execute the steps of the interactive version of the proof system in an incorrect order, likely breaking the assumptions in the security proof. In particular, the issue enables a malicious prover, to create an invalid proof of knowledge for some public input given a valid proof for another public input.
A.3 Soundness and Malleability in Nova IVC
@@
In Nova IVC at each new round the statement regarding the prefix of computation so far is folded with the statement regarding the final step, this folding has to be done in circuit which could involve non-native arithmetic operations that are relatively costly. One approach to deal with non-native arithmetic operations is to use cycles of elliptic curves, switching the circuit and curve at each step, so that almost all of the in-circuit arithmetic operations are native. The original implementation of Nova indeed used a 2-cycle of curves. This variation of the proof system was not fully specified nor proved to be secure in the original paper. Nguyen et al. [70] showed that the implementation misses a critical check to verify the consistency of the inputs to the two circuits, breaking the soundness completely. Notably, [70] proposes a fix for this issue and provides proof of security.

### v4

@@ 1 Introduction
On the forefront of the practical application of general-purpose ZKPs are Succinct Non-interactive Arguments of Knowledge (SNARKs) [50, 87, 56, 46, 29]. SNARKs are non-interactive protocols that allow the prover to generate a succinct proof. The proof is efficiently checked by the verifier, while maintaining three crucial properties: completeness, soundness, and zero-knowledge. What makes SNARKs particularly appealing is their general-purpose nature, allowing any computational statement represented as a circuit to be proven and efficiently verified. Typically, SNARKs are used to prove that for a given function $f$ and a public input $x$, the prover knows a (private) witness $w$, such as $f(x,w)=y$. This capability allows SNARKs to be used in various applications, including ensuring data storage integrity [94], enhancing privacy in digital asset transfers [75, 97] and program execution [18, 16], as well as scaling blockchain infrastructure [91, 90, 67, 99]. Their versatility also extends to non-blockchain uses, such as in secure communication protocols [110, 69] and in efforts to combat disinformation [61, 63]. Unfortunately, developing and deploying systems that use SNARKs safely is a challenging task.
@@ 2 Background on SNARKs
$\textit{Setup}(pp)\rightarrow(pk,vk)$. Given public parameters $pp$ as input, output proving and verification keys $pk$ and $vk$.
$\textit{Prove}(pk,x,w)\rightarrow\pi$. Given the proving key $pk$, the public input $x$, and the witness $w$, output a proof $\pi$.
$\textit{Verify}(vk,x,\pi)\rightarrow\{0,1\}$. Given the verification key $vk$, the public input $x$, and the proof $\pi$, output $1$ if the proof is valid and 00 otherwise.
@@ 3.1 System Model
Circuit Embedded Domain Specific Languages. An Embedded Domain Specific Language (eDSL) is a type of domain-specific language that is embedded within a host general-purpose programming language. For developing SNARKs, several eDSLs have emerged in recent years, embedded as libraries in Golang [17], Rust [109, 41, 93], and TypeScript [78]. eDSLs can seamlessly interact with other code written in the host language, allowing easy integration with existing libraries. At the same time, they require developers to actively distinguish between in-circuit and out-of-circuit operations, which requires domain expertise to ensure correct implementation.
@@
Circuit Compiler. The circuit compiler in a SNARK implementation compiles the circuit specified in a high-level DSL or eDSL to the respective arithmetization. Note, that not all compilers target the compilation of high-level languages to specific arithmetizations. For example, high-level programs can first be compiled to VampIR  [^1] https://github.com/anoma/vamp-ir and ACIR  [^2] https://github.com/noir-lang/acvm, which are intermediate representations that allow compilation to, e.g., R1CS or some form of Plonkish arithmetization and aim to ease the support of multiple backends. Similarly, zkLLVM provides a compiler from the LLVM IR to different arithmetizations, allowing users to prove the execution of, e.g., native Rust, Golang, or C++  [^3] https://github.com/NilFoundation/zkLLVM.
@@
Setup. In pre-processing SNARKs, the setup algorithm generates the prover and verifier keys ($pk$ and $vk$) by encoding the circuit relation, with the primary goal of succinct verification and optimized proving. However, different SNARKs have vastly differing properties with regard to trust assumptions, which in turn also impact their performance. Some SNARKs require a trusted setup, i.e., a randomized setup phase that introduces a trapdoor either per circuit (i.e., non-universal) [56] or in a universal setup phase [46]. The randomized part of the setup involves public parameters and random input to generate a common reference string (CRS) [77], which can be used to derive $pk$ and $vk$ in the deterministic part of the setup for a concrete circuit instantiation. The trapdoor needs to be discarded after the setup proccess is done, as anyone who has knowledge of the trapdoor can forge proofs. SNARKs that do not employ a trapdoor in the setup phase are commonly coined transparent [13]. Note, that this introduces a trade-off in verifier efficiency — transparent SNARKs only have poly-logarithmic-time verification as compared to constant-time for SNARKs with trusted setup [102].
@@
Prover & Verifier. The prover consumes the values generated by the frontend, where the circuit is represented in a specific arithmetization. At a high level, the prover of a pre-processing SNARK (aside from Groth16 [56]) utilizes a polynomial commitment to commit to the satisfying witness, and successively engages with the verifier in an interactive protocol to evaluate the correctness of assignments with regard to the circuit relation. This “recipe” leads to a SNARK if the verifier is public-coin [88], and the prover can hash the transcript of the interactive proof to render the protocol non-interactive (a process commonly denoted as applying the “Fiat-Shamir heuristic”). Different proof systems choose different combinations of polynomial commitments and interactive proofs, resulting in differing properties for the respective target applications. The most popular approaches include Plonk [46], which employs a constant-round polynomial interactive oracle proof (IOP) with the KZG commitment [62]. Halo2 [109] uses the Plonk IOP with the Bulletproofs polynomial commitment[20]. Another class of protocol combines a polynomial commitment based on Merkle hashes with a low-degree test, resulting in a protocol that does not demand for a trusted setup or operations in elliptic curve groups [12], at the cost of non-constant verification time and a larger proof size [39]. Plonky2 [93] uses the FRI-based polynomial commitment scheme and combines it with the Plonk IOP. Groth16 [56], the SNARK with the shortest proof size, is based on probabilistically checkable proofs (PCP) [50] and pairings over elliptic curves for proof verification. For an in-depth taxonomy, we refer readers to [102]. Some applications leverage one of the above SNARKs without the zero-knowledge property, solely relying on non-interactivity and succinctness (e.g zk-EVMs [99, 91]). In most SNARKs, zero-knowledge can be obtained cheaply with minor modifications.
@@ 4 Methodology
Analyzed SNARK Implementations. Our examination focuses on widely deployed SNARK systems, including ZK-rollups for blockchain scalability (e.g., zkSync Era [67], Polygon ZK-EVM [91] Scroll [99]), [^4] Notably, ZK-rollups have more than 1B USD in accumulated value locked in them according to L2BEAT https://l2beat.com/scaling/summary privacy-centric blockchains (e.g., Zcash [97], Aztec [6], and decentralized privacy applications (e.g., Tornado Cash [89]). We also assess key circuit libraries and SNARK frameworks (e.g., Circom and halo2) [39]. Our analysis further extends to inspecting the underlying proof systems for known bugs. This comprehensive approach ensures that we cover a broad spectrum of real-world SNARK applications and the various layers of technology they rely on.
Data Sources. In our comprehensive study of security vulnerabilities related to SNARK systems, we have utilized an extensive range of data sources to ensure a thorough analysis. Our primary source includes critical and high-impact vulnerabilities identified in $107$ security audit reports of systems employing SNARKs. Among these, $75$ reports specifically highlighted vulnerabilities directly related to SNARKs, while the remaining reports detailed other types of vulnerabilities, such as common smart contract deficits. Additionally, we have incorporated vulnerabilities from public bug bounty programs and vulnerability disclosures. Our search for disclosures included those from prominent projects like Zcash, Aztec Connect, and TornadoCash, disclosures by security firms and researchers specializing in ZKPs, and comprehensive web searches targeting ZKP vulnerabilities. We reviewed Web3 bug bounty programs like Immunefi, though we found no ZKP-related disclosures there. Additionally, we consulted closely with top audit firms focusing on ZKPs to ensure we didn’t miss any critical disclosures. These sources provide real-world examples of how vulnerabilities have impacted operational systems as well as the measures taken to address them. Furthermore, we delved into bug trackers of popular GitHub projects related to SNARKs. Our focus here has been specifically on security-related bugs, filtering out non-security-related issues. Table 2 in our paper depicts these varied data sources, collectively forming the foundation of our study of SNARK security.
@@ 8 Defenses For SNARKs
Despite these initial strides in tool development for ZK circuits, significant limitations remain. Tools relying on SMT solvers, like Picus and Korrect, face challenges due to limited support and efficiency in handling finite field arithmetic, leading to performance bottlenecks [58, 83]. Static analysis tools are often restricted in the range of vulnerabilities they can detect and tend to be specific to certain languages. CirC [82] compiler represents a step towards language-agnostic tooling, allowing various DSLs to compile into an Intermediate Representation, which can then be analyzed for potential vulnerabilities. Another promising direction is the creation of more secure DSLs, like Aleo [31], noname  [^8] https://github.com/zksecurity/noname, and Noir  [^9] https://aztec.network/aztec-nr/, or eDSLs such as o1js [78], which incorporate strong typing and improved abstractions to prevent common issues.
@@
Multi-Provers as a Defense Mechanism. SNARK proofs are often used in safety-critical contexts, yet, with the current state of technology, guaranteeing that a complex SNARK system is bug-free and safe remains an elusive goal. The multi-prover design enhances safety by introducing redundancy at the proof system level. A multi-prover system utilizes multiple proof systems besides the primary SNARK, such as alternative SNARK proofs or trusted-execution environment (TEE) based proofs [22], and primary proofs are accepted only if the secondary ones agree with it. This design trades off liveness with more safety, as a dispute will potentially bring the system to a halt. Such systems require a dispute resolution mechanism to handle potential disagreements. In practice, some ZK-Rollup projects [99], arguably one of the most complex categories of SNARK systems today, are already experimenting with multi-provers. [^10] https://scroll.io/blog/scaling-security
@@ 9 Issues in Proof Systems
Universal Composability. The universal composability (UC) framework by Canetti [21] is widely considered as the “gold standard” for proving the security of cryptographic primitives. The reason for the popularity of the UC model is that it can guarantee strong security against adaptive adversaries and further allows for modular reusability of cryptographic primitives in greater, high-level protocol designs. Arguing UC security for a SNARK, or a SNARK-based protocol, is non-trivial, as SNARKs commonly use techniques that are not realizable in the UC model [55], and result in SNARKs that are not formally non-malleable. Recent work aims to close this gap by studying compilers that render common SNARKs UC-secure by encrypting the witness and including it in the argument, which results in an overhead that renders the SNARK non-succinct due to the unbounded size of the witness [65, 1, 2, 7]. Most recent work proves that initial SNARK constructions [74, 14] that do utilize random oracles are UC-secure [28].
In practice, we observe that UC security is often not considered, and non-UC-secure protocols, like Groth16 [56], are applied without further ado. However, we are not aware of any protocol that claims UC-security and successively got exploited due to erroneous analysis, i.e., an underspecified ideal functionality or incorrectly claiming that the real-life model can be simulated in the ideal process model.
@@ 10 Discussion
RQ: What tools and methodologies can be developed to better identify and mitigate under-constrained bugs in ZK circuits? Which techniques are the most efficient in detecting such bugs?
2) Insight – Soundness bugs affecting SNARK verifiers lead on average to high severity bugs. Particularly when soundness bugs are exploitable, typically due to under-constrained circuits, they can lead to significant security breaches. Their potential to compromise system security has led to suboptimal strategies, such as employing multi-provers and permissioned provers, alongside significant bug bounties, sometimes reaching up to \$500k for a single vulnerability.  [^11] https://aztec.network/blog/aztec-connect-postmortem/ These measures underscore the critical nature of soundness in maintaining the trustworthiness of SNARK-based systems.
RQ: What are effective strategies to detect and prevent soundness bugs in SNARKs? Can we formally verify the verifiers, as their scope is limited?
3) Insight – Low-level circuit DSLs are easy-to-misuse leading to many vulnerabilities. Crafting efficient ZK circuits often necessitates using low-level ZK DSLs such as Circom and Gnark, reminiscent of early assembly and C programming, where common high-level programming features such as abstractions and basic types are absent. This complexity not only steepens the learning curve for developers but also increases the likelihood of introducing vulnerabilities into the circuits. The prevalent use of such DSLs highlights a regression to an era with bugs due to the absence of modern programming safeguards, underscoring an urgent need for more user-friendly DSLs that offer better abstractions and safety features, thereby mitigating the main shortcomings and reducing the vulnerability surface of SNARK applications.
RQ: How can more user-friendly DSLs be designed to reduce the vulnerability surface of SNARK applications?
@@
RQ: What educational and tooling resources can assist developers and auditors in navigating the complexities of using SNARKs? How useful can specifications (formal or informal) be towards detecting vulnerabilities in systems using ZKPs?
@@
RQ: How can the reliability and security of compilers and proof system implementations for SNARKs be ensured? Which testing or verification techniques can be applied?
@@
RQ: What improvements are needed in security tools to effectively scale and cover a broader range of vulnerabilities? What are the limitations of formal verification tools?
@@
RQ: What is the security of SNARKs instantiation given specific curves and fields? How many bits of security does each SNARK have when using specific configurations?
@@
RQ: How can we develop detection mechanisms or heuristics for exploits in privacy-preserving SNARK systems?
@@ Acknowledgments
We would like to thank our shepherd and the anonymous USENIX reviewers. We also thank Shankara Pailoor, Alex Ozdemir, Jean-Philippe Aumasson, Oana Ciobotaru, and Justin Traglia for their insightful feedback. Finally, we thank the Ethereum Foundation and 0xPARC for supporting this work.
@@ A.2 Frozen Heart – Insecure Fiat Shamir Transformation
Note that if at any point a malicious prover does not like the challenge produced by Fiat-Shamir, his only option is to change his previous messages, but given that the hash function is cryptographically secure, he is unable to efficiently find a suitable change. However, invoking Fiat-Shamir with only a partial transcript of the protocol can create serious soundness issues, referred to as Frozen Heart attacks [36, 80]. Intuitively, this enables a malicious prover to execute the steps of the interactive version of the proof system in an incorrect order, likely breaking the assumptions in the security proof. In particular, the issue enables a malicious prover, to create an invalid proof of knowledge for some public input given a valid proof for another public input.
A.3 The Last Challenge Attack
The Last Challenge Attack [32] specifically exploits a flaw in KZG-based SNARK implementations where the final FS transform challenge (i.e., the KZG batching challenge) is computed by a non-compliant verifier using only a part of the full communication transcript. In this scenario, a malicious prover can manipulate the proof process to produce a valid proof for a false statement. The attack occurs because the malicious prover can set to arbitrary values of its choice all public inputs as well as some of the components of the SNARK proof, and for the rest of the components, the prover can, with high probability, compute values that pass the non-compliant verification check. Consequently, the malicious prover can create an invalid proof for some public input by leveraging a valid proof for another public input. This can lead to disastrous consequences in practice.
While both the Last Challenge Attack and the Frozen Heart vulnerability exploit incorrect implementations of the FS transform, and both may allow for creating proofs of incorrect statements, they differ in specifics. The Frozen Heart vulnerability may occur if parts of the public input are omitted by the FS transform. The Last Challenge Attack may occur if one omits parts of the transcript, other than the public input, in the computation of the final FS transform challenge.
A.4 Soundness and Malleability in Nova IVC
@@
In Nova IVC, at each new round, the statement regarding the prefix of computation so far is folded with the statement regarding the final step; this folding has to be done in the circuit, which could involve non-native arithmetic operations that are relatively costly. One approach to deal with non-native arithmetic operations is to use cycles of elliptic curves, switching the circuit and curve at each step, so that almost all of the in-circuit arithmetic operations are native. The original implementation of Nova indeed used a 2-cycle of curves. This variation of the proof system was not fully specified nor proved to be secure in the original paper. Nguyen et al. [76] showed that the implementation misses a critical check to verify the consistency of the inputs to the two circuits, breaking the soundness completely. Notably, [76] proposes a fix for this issue and provides proof of security.